A certain reaction has an activation energy of 60.0 kJ/mol and a frequency factor of A1 = 2.10×1012 M−1s−1 . What is the rate constant, k, of this reaction at 29.0 ∘C ?

Express your answer with the appropriate units. Indicate the multiplication of units explicitly either with a multiplication dot (asterisk) or a dash.
Answer #1

Activation Energy (Ea) = 6010 ks/mol = 60 x103 g/mol frequency factor (A,)= 2-10 x102 Msec- Temperature (7) = 29% = (29+213)